# spacing=nonfrench

### v1

@@ 4.2. Cache Admission and Eviction Workflows
Data structures. FMCache monitors path access frequencies for read operations (excluding access to ancestors during path resolution). For uncached paths, which dominate path access traffic, FMCache uses a Count-Min Sketch (CMS) (Cormode and Muthukrishnan, 2005), as in prior studies (Jin et al., 2017; Liu et al., 2019; Sheng et al., 2023), to estimate their access frequencies within fixed-size memory with provable error bounds. For cached paths, FMCache tracks their exact access frequencies using a frequency counter array. The switch periodically reports the access frequencies of all cached paths to the controller for eviction decisions, and resets the CMS and frequency counter array after each reporting.
@@
Cache admission. FMCache’s cache admission is triggered by the switch data plane. During runtime, the switch monitors the access frequencies of uncached paths via the CMS and identifies hot paths (i.e., those exceeding a pre-defined CMS threshold) for admission. When a hot path $p$ is detected, the switch notifies the controller, which retrieves the metadata for $p$ and its uncached ancestors from the servers. The controller communicates with servers using UDP (and issues retransmissions if needed), bypassing the switch data plane to avoid critical-path overhead, as in prior work (Jin et al., 2017; Sheng et al., 2023).
@@
The controller verifies cache capacity. If the cache is not full, the controller sends the hash keys and metadata for $p$ and its uncached ancestors to the switch for admission (note that the controller maintains a global view of all cached paths). Unlike NetCache (Jin et al., 2017) (without path awareness), FMCache admits the metadata for both $p$ and its uncached ancestors to ensure path-aware caching.
@@
Cache eviction. When the cache is full and an uncached hot path $p$ is reported, the controller triggers cache eviction to reclaim cache space for $p$ and its uncached ancestors. It selects cached path candidates for eviction based on periodically reported access frequencies, prioritizing the least frequently accessed path with no cached descendants. If the selected path is the only cached child of its parent, the parent is also selected. The controller recursively includes ancestors as candidates until reaching an ancestor with multiple cached children or the root. The rationale of this recursive inclusion is that the selected ancestors are unlikely accessed by themselves, as directory operations are infrequent in practice (Ren et al., 2014; Lv et al., 2022; Xu et al., 2024) (e.g., only 4.2% of metadata operations in Alibaba’s Pangu file system are directory-related (Lv et al., 2022)).
@@
Figure 3. Example of cache admission and eviction workflows.
@@
5. In-Switch Read-Write Locking
@@
FMCache adopts multi-level read-write locking to ensure cache consistency under concurrent metadata updates. The locking mechanism operates entirely within the switch data plane to avoid controller overhead.
5.1. Lock Design
@@
FMCache employs multiple lock counter arrays and a validation array, both implemented as register arrays in the switch data plane and initialized with zero entries. Each cached path is associated with one lock counter (a slot in a lock counter array) and one validation flag (a slot in the validation array), indexed by the path’s hash key.
@@
Lock counter arrays. Each lock counter array corresponds to specific path levels. Each of its lock counters records the number of active read requests for a cached path at that level. Due to switch resource constraints, FMCache allocates eight counter arrays of 65,536 16-bit counters each, while each counter supports up to 65,535 concurrent read requests. The first seven arrays are assigned to levels 1 through 7 (e.g., /a for level 1 (§2.1)), while the eighth array handles all remaining deeper levels (using the hash key of level 8). This design maximizes concurrent access to shallow paths, based on empirical evidence that most metadata requests are aggregated at small depths (Agrawal et al., 2007; Meyer and Bolosky, 2012; Douceur and Bolosky, 1999); for example, 90% (Agrawal et al., 2007; Meyer and Bolosky, 2012), or even close to 100% (Douceur and Bolosky, 1999), of accessed paths have a depth of no more than 10. FMCache maps a cached path to a lock counter array based on its level, and uses its hash key’s last 16 bits to associate the path with a specific lock counter.
@@
Validation array. The validation array (Jin et al., 2017; Sheng et al., 2023) tracks metadata validity for all cached paths. A validation flag of one indicates valid metadata and allows reads from the cache, while a flag of zero indicates invalid or updating metadata and all reads are directed to servers. After a cache update, its validation flag is set to one to permit subsequent reads.
5.2. Read and Write Flows
@@
FMCache places the validation array across all egress pipelines, co-located with value register arrays (§4.1) for efficient validity checks. Lock counter arrays reside in a single ingress pipeline to avoid cross-pipeline synchronization. FMCache redirects requests arriving at other ingress pipelines to the ingress pipeline holding lock counter arrays via cross-pipeline recirculation (§7). This placement is critical, as lock counter arrays should be positioned before the validation array to ensure correctness, and they cannot be placed in egress pipelines due to insufficient switch resources. In FMCache’s deployment, the switch data plane is not a bottleneck, and consistent packet processing ordering is maintained across ingress and egress pipelines, so FMCache ensures efficient and correct locking and validation operations.
@@
Reads. Reads are classified as (i) single-path reads (e.g., stat), which retrieve only metadata of the requested path; and (ii) multi-path reads (e.g., readdir), which retrieve metadata of the requested path and its descendants. Single-path reads are served from the in-switch cache, while multi-path reads are forwarded to servers to ensure correctness, as descendant paths may be uncached and servers maintain the authoritative namespace to resolve partially caching scenarios. Multi-path reads are rare in practice (e.g., only 3.9% in Alibaba’s workloads (Lv et al., 2022); see Table 1). Upon receiving a (single-path) read request, the switch checks if the path’s last level is cached, implying that all its ancestors are also cached (§4). If so, the switch increments the lock counter for each path level by one and resolves the path via recirculation.
@@
Any ACK loss can cause duplicate lock decrements via retransmissions and violate correctness. FMCache handles ACK loss via a server-switch sequence-number protocol. Each server tags a lock-related response with a local sequence number (initially 0), and increments the sequence number upon receiving an ACK from the switch. The switch also maintains per-server expected sequence numbers in a sequence counter array. Upon receiving a server’s response, the switch compares the embedded sequence number with its expected value: if they match, the switch increments the expected value (for the next response), processes lock updates, forwards the response to the client, and returns an ACK to the server; if the embedded sequence number is lower, it indicates a duplicate, so the switch only sends an ACK to the server to suppress further retransmissions.
Figure 4. Example of processing a read request under multi-level read-write locking.
@@
Note that a write delayed by ongoing reads may be overtaken by a new read arriving before recirculation, so it is theoretically possible for a write to be blocked indefinitely by continuous reads. However, the short recirculation time makes this unlikely. FMCache’s locking ensures correctness and allows the write to proceed after sufficient recirculations.
@@
(ii) Multi-path writes. Multi-path writes follow the same single-path write processing until the request reaches the server. If the write is successfully completed, the server updates the cache for all cached descendants before the requested path, so that the requested path remains invalidated until all cached descendants are fully updated. By performing path resolution in a top-down manner, FMCache prevents reads from accessing cached descendants before the requested path is updated (i.e., until the cache updates for a multi-path write are completed), so as to ensure cache consistency.
6. Local Hash Collision Resolution
@@
Figure 5. Example of token generation and distribution. Note that the controller also assigns tokens for $p$’s uncached ancestors in Steps 2-4. We omit the details for brevity.
@@
FMCache maps file-system pathnames to 64-bit hash keys (§4.1), so hash collisions are possible (albeit unlikely) and lead to incorrect metadata retrieval. While the controller holds a global view of all cached paths (§4.2), querying the controller to resolve hash collisions for every request is impractical due to high switch-to-controller latencies (§2.3). FMCache adopts a local, token-based hash collision resolution mechanism by synchronizing the controller’s global view of cached paths with the switch, clients, and servers, so as to ensure correctness without compromising performance.
@@
Hash collision resolution with tokens. A token is an 8-bit value paired with a 64-bit hash key to uniquely identify a path. Valid tokens range from 1 to 255, while 0 indicates invalid. A cached path is assigned token 1 if no collision occurs, or the next available token (e.g., 2) if it collides with an existing cached path. Thus, we can associate valid tokens with up to 255 paths with the same colliding hash key. The probability of two paths colliding is $1/{2}^{32}$ based on the birthday paradox (Flajolet et al., 1992), making it unlikely for more than 255 paths to share the same hash key in a large namespace. Thus, 8-bit tokens suffice for hash collision resolution.
@@
FMCache maintains two unordered map structures: (i) a path-token map, which records paths and their tokens (path-token pairs), and (ii) a hash-token map, which records hash keys and their tokens (hash-token pairs). The controller keeps both maps, each client and server holds a path-token map, and the switch holds a hash-token map.
@@
Token generation and distribution. During cache admission, the controller assigns tokens to each level of a hot path, as shown in Figure 5. When the switch reports a hot path $p$ for admission ( 1), the controller checks its map structures. If $p$ is the first time being admitted and its hash key is unique, the controller assigns a token of value one; if a collision occurs, it assigns the next available token (e.g., $t$) ( 2). The controller updates its path-token and hash-token maps with the new path-token and hash-token pairs, respectively ( 3). These entries persist in the controller’s maps even after cache eviction to ensure consistency. If an evicted path is later re-admitted, its previously assigned token is reused. In the worst case, the controller may store entries for all paths ever cached. For scalability, the controller can use persistent key-value stores (e.g., RocksDB (roc, [n. d.])) for the map structures.
@@
Figure 6. Example of how a client updates its path-token map. Note that the client also attaches tokens of $p$’s ancestors in Step 1, the server returns the tokens for $p$’s ancestors in Step 5, and the client adds the tokens in Step 6. We omit the details for brevity.
@@
Discussion. FMCache’s token-based mechanism ensures correctness with modest client complexity and overhead. Clients must maintain path-token maps and attach tokens to requests, and incur a cache miss on the first read of a cached path due to the initial lack of a valid token. Also, including tokens in requests adds minor communication overhead. Theoretically, the number of tokens grows linearly with the depth of a path. Nevertheless, as most paths have small depths (§5.1), the communication overhead is minimal.
7. Implementation
We have built a prototype of FMCache, including the controller, in-switch cache, clients, and servers.
Controller. The controller is implemented in C++ with 2.8 K LoC. It leverages APIs provided by the Tofino switch compiler (tof, [n. d.]) to interact with the switch data plane, including configurations and updates of MATs and registers, for cache admission and eviction (§4) and token management (§6).
In-switch cache. The in-switch cache is implemented in P4 with 5.1 K LoC and compiled for a Tofino switch (tof, [n. d.]; swi, [n. d.]). It comprises several components: (i) a hash-token map using MATs for cache lookups and local hash collision resolution, where each entry is a 9-byte key (8-byte hash key and 1-byte token); (ii) 32 register arrays of 32-bit slots for storing cache entries; (iii) a three-row CMS with three register arrays, each with 64K 16-bit slots; (iv) a frequency counter array with a register array of 32-bit slots; (v) lock counter arrays with eight register arrays, each with 16K 16-bit slots; (vi) a validation array with a register array of 1-bit slots; and (vii) a sequence counter array with a register array of 8-bit slots.
Client implementation. The C++ client driver supports multi-threaded workload execution, where each thread simulates a logical client. Each client maintains a path-token map for local hash collision resolution (§6). It communicates with the switch via UDP-encapsulated packets customized for metadata requests, with re-transmission support for reliability. We integrate the driver with the mdtest (mdt, [n. d.]) benchmarking tool for HDFS metadata operations.
Server implementation. Each server is implemented in C++ with 3 K LoC and hosts an HDFS (v3.2) namenode (hdf, [n. d.]a). We use RBF (hdf, [n. d.]b) with the HASH_ALL policy, which uses consistent hashing to distribute files evenly across all namenodes for load balancing and creates directories on all namenodes. Each server manages a subset of the metadata namespace, connects to its local namenode via the C++ HDFS client library libhdfs3 (lib, [n. d.]), and serves client requests while maintaining a path-token map for local hash resolution (§6).
8. Evaluation
8.1. Methodology
@@
Alibaba (Lv et al., 2022) 52.6% open/close, 9.59% create, 3.9% readdir, 69.1%
@@
Training (Xu et al., 2024) 54.32% open/close, 28.5% stat, 0.13% readdir, 83.1%
@@
Thumb (Xu et al., 2024) 57.01% open/close, 28.44% stat, 0.13% readdir, 85.7%
@@
LinkedIn (Ren et al., 2014) 84% open/getattr, 9% create/mkdir, 84%
@@
Table 1. Real-world workloads and their metadata operation ratios.
@@
Testbed. We evaluate FMCache on a testbed comprising a 3.2 Tbps two-pipeline Tofino switch (tof, [n. d.]; swi, [n. d.]) and three physical machines. Two machines host servers, each with a 2.40 GHz 10-core Intel Xeon Silver 4210R CPU, 128 GiB DRAM, and a 2 TB HDD (Dell PERC H330 Mini). The client driver runs on the remaining machine, with a 2.40 GHz, 16-core Intel Xeon Silver 4314 CPU, 128 GiB DRAM, and a 960 GB NVMe SSD (Micron 9300 PRO). Each machine is connected to the switch via a 40 Gbps NIC (Mellanox ConnectX-5 CX516A). The client machine uses one pipeline, and the two server machines use another pipeline. All counter arrays (§7) reside in the server-connected pipeline.
@@
Since our two-pipeline Tofino switch lacks native support for cross-pipeline recirculation, we physically connect the designated ingress pipeline hosting the lock counter arrays to another ingress pipeline using a physical wire (Sheng et al., 2023). This enables requests from another ingress pipeline to be recirculated to the designated ingress pipeline (§5). Future programmable switches with native cross-pipeline recirculation would eliminate this requirement.
@@
Workloads. We evaluate FMCache using four real-world workload traces: (i) Alibaba’s Pangu file-system instances (Alibaba) (Lv et al., 2022), (ii) convolutional neural network training (Training) (Xu et al., 2024), (iii) the processing of one million thumbnail images (Thumb) (Xu et al., 2024), and (iv) a LinkedIn HDFS cluster (LinkedIn) (Ren et al., 2014). Table 1 summarizes the proportion of metadata operations and the read ratio for each workload.
@@
We refine the workloads for our evaluation. To focus on metadata performance, we exclude reads and writes of file data, following prior studies (Niazi et al., 2017; Li et al., 2017a; Ren et al., 2014; Lv et al., 2022). For Training and Thumb that include file reads and writes (Xu et al., 2024), we exclude these operations and normalize the remaining metadata operations. Since we exclude file writes, we treat close as a read operation, while it updates both modification and access timestamps if the closed file has been updated. HDFS updates access timestamps hourly by default (def, [n. d.]), and we exclude timestamp updates as they are infrequent.
@@
For LinkedIn, as the original paper (Ren et al., 2014) does not provide operation ratios, we adjust the ratios based on trace analysis (Lv et al., 2022) as follows: open (42%), getattr (42%), create (4.5%), mkdir (4.5%), chmod (1%), delete (3%), and rename (3%). We assign the smallest ratio to chmod as it is less frequent than delete and rename (Lv et al., 2022). Further, getattr corresponds to stat and statdir in HDFS. We replace getattr with stat, as file operations are much more frequent than directory operations (Lv et al., 2022), and focus on file renaming for rename.
@@
Generation of metadata operations. We use mdtest (mdt, [n. d.]) to generate file-system namespaces and metadata operations. By default, we configure a path depth of nine, as metadata requests are often aggregated at small depths (Agrawal et al., 2007; Meyer and Bolosky, 2012; Douceur and Bolosky, 1999)5.1). We create 32 million empty files to focus on metadata performance; empty files are also used in prior evaluation (Ren et al., 2014; Xiao et al., 2015; Niazi et al., 2017; Li et al., 2017a; Lv et al., 2022; Wang et al., 2023). To simulate workload skewness, we use the 80/20 rule, with 80% of operations on 20% of directories and files (Wang et al., 2021a; Xu et al., 2024). We model file access frequencies across levels using a power-law distribution with an exponent of 0.9 and randomly assigning an operation type using the ratios in Table 1. For statdir and readdir, we choose the parent directory of the selected file. For mkdir and rmdir, we use separate directories to avoid removing non-empty directories. We address various workload settings in §8.3.
@@
We mix all metadata operations in each workload. However, rename, delete, and rmdir inherently involve metadata modifications that necessitate the granting and revoking of leases in HDFS (Wang et al., 2009). The lease-based operations on frequently created and deleted files can slow down all metadata operations when we perform stress tests. Thus, we place rename, delete, and rmdir at the end of the request sequence, so as to keep metadata operations at high rate.

### v2

@@ IV-B Cache Admission and Eviction Workflows
Data structures. Fletch monitors path access frequencies for read operations (excluding access to ancestors during path resolution). For uncached paths, which dominate path access traffic, Fletch uses a Count-Min Sketch (CMS) [cormode05], as in prior studies [jin17, liu19, sheng25], to estimate their access frequencies within fixed-size memory with provable error bounds. For cached paths, Fletch tracks their exact access frequencies using a frequency counter array. The switch periodically reports the access frequencies of all cached paths to the controller for eviction decisions, and resets the CMS and frequency counter array after each reporting.
@@
Cache admission. Fletch’s cache admission is triggered by the switch data plane. During runtime, the switch monitors the access frequencies of uncached paths via the CMS and identifies hot paths (i.e., those exceeding a pre-defined CMS threshold) for admission. When a hot path $p$ is detected, the switch notifies the controller, which retrieves the metadata for $p$ and its uncached ancestors from the servers. The controller communicates with servers using UDP (and issues retransmissions if needed), bypassing the switch data plane to avoid critical-path overhead, as in prior work [jin17, sheng25].
@@
The controller verifies cache capacity. If the cache is not full, the controller sends the hash keys and metadata for $p$ and its uncached ancestors to the switch for admission (note that the controller maintains a global view of all cached paths). Unlike NetCache [jin17] (without path awareness), Fletch admits the metadata for both $p$ and its uncached ancestors to ensure path-aware caching. To guarantee cache coherence during cache admission, while the controller is adding a path to the cache, write requests to this path are blocked at the server until the server is notified that the cache admission is finished [jin17]. Since file-system metadata workloads are read-intensive (§III-A), the blocking action has limited impact on the overall performance in general.
@@
Cache eviction. When the cache is full and an uncached hot path $p$ is reported, the controller triggers cache eviction to reclaim cache space for $p$ and its uncached ancestors. It selects cached path candidates for eviction based on periodically reported access frequencies, prioritizing the least frequently accessed path with no cached descendants. If the selected path is the only cached child of its parent, the parent is also selected. The controller recursively includes ancestors as candidates until reaching an ancestor with multiple cached children or the root. The rationale of this recursive inclusion is that the selected ancestors are unlikely accessed by themselves, as directory operations are infrequent in practice [ren14, lv22, xu24] (e.g., only 4.2% of metadata operations in Alibaba’s Pangu file system are directory-related [lv22]).
@@
Figure 3: Example of cache admission and eviction workflows.
@@
V In-Switch Read-Write Locking
@@
Fletch adopts multi-level read-write locking to ensure cache consistency under concurrent metadata updates. The locking mechanism operates entirely within the switch data plane to avoid controller overhead.
V-A Lock Design
@@
Fletch employs multiple lock counter arrays and a validation array, both implemented as register arrays in the switch data plane and initialized with zero entries. Each cached path is associated with one lock counter (a slot in a lock counter array) and one validation flag (a slot in the validation array), indexed by the path’s hash key.
@@
Lock counter arrays. Each lock counter array corresponds to specific path levels. Each of its lock counters records the number of active read requests for a cached path at that level. Due to switch resource constraints, Fletch allocates eight counter arrays of 65,536 16-bit counters each, while each counter supports up to 65,535 concurrent read requests. The first seven arrays are assigned to levels 1 through 7 (e.g., /a for level 1 (§II-A)), while the eighth array handles all remaining deeper levels (using the hash key of level 8). This design maximizes concurrent access to shallow paths, based on empirical evidence that most metadata requests are aggregated at small depths [agrawal07, meyer12, douceur99]; for example, 90% [agrawal07, meyer12], or even close to 100% [douceur99], of accessed paths have a depth of no more than 10. Fletch maps a cached path to a lock counter array based on its level, and uses its hash key’s last 16 bits to associate the path with a specific lock counter.
@@
Validation array. The validation array [jin17, sheng25] tracks metadata validity for all cached paths. A validation flag of one indicates valid metadata and allows reads from the cache, while a flag of zero indicates invalid or updating metadata and all reads are directed to servers. After a cache update, its validation flag is set to one to permit subsequent reads.
V-B Read and Write Flows
@@
Fletch places the validation array across all egress pipelines, co-located with value register arrays (§IV-A) for efficient validity checks. Lock counter arrays reside in a single ingress pipeline to avoid cross-pipeline synchronization. Fletch redirects requests arriving at other ingress pipelines to the ingress pipeline holding lock counter arrays via cross-pipeline recirculation (§VII). This placement is critical, as lock counter arrays should be positioned before the validation array to ensure correctness, and they cannot be placed in egress pipelines due to insufficient switch resources. In Fletch’s deployment, the switch data plane is not a bottleneck, and consistent packet processing ordering is maintained across ingress and egress pipelines, so Fletch ensures efficient and correct locking and validation operations.
@@
Reads. Reads are classified as (i) single-path reads (e.g., stat), which retrieve only metadata of the requested path; and (ii) multi-path reads (e.g., readdir), which retrieve metadata of the requested path and its descendants. Single-path reads are served from the in-switch cache, while multi-path reads are forwarded to servers to ensure correctness, as descendant paths may be uncached and servers maintain the authoritative namespace to resolve partially caching scenarios. Multi-path reads are rare in practice (e.g., only 3.9% in Alibaba’s workloads [lv22]; see Table I). Upon receiving a (single-path) read request, the switch checks if the path’s last level is cached, implying that all its ancestors are also cached (§IV). If so, the switch increments the lock counter for each path level by one and resolves the path via recirculation.
@@
Any ACK loss can cause duplicate lock decrements via retransmissions and violate correctness. Fletch handles ACK loss via a server-switch sequence-number protocol. Each server tags a lock-related response with a local sequence number (initially 0), and increments the sequence number upon receiving an ACK from the switch. The switch also maintains per-server expected sequence numbers in a sequence counter array. Upon receiving a server’s response, the switch compares the embedded sequence number with its expected value: if they match, the switch increments the expected value (for the next response), processes lock updates, forwards the response to the client, and returns an ACK to the server; if the embedded sequence number is lower, it indicates a duplicate, so the switch only sends an ACK to the server to suppress further retransmissions.
Figure 4: Example of processing a read request under multi-level read-write locking.
@@
Note that a write delayed by ongoing reads may be overtaken by a new read arriving before recirculation, so it is theoretically possible for a write to be blocked indefinitely by continuous reads. However, the short recirculation time makes this unlikely. Fletch’s locking ensures correctness and allows the write to proceed after sufficient recirculations.
@@
(ii) Multi-path writes. Multi-path writes follow the same single-path write processing until the request reaches the server. If the write is successfully completed, the server updates the cache for all cached descendants before the requested path, so that the requested path remains invalidated until all cached descendants are fully updated. By performing path resolution in a top-down manner, Fletch prevents reads from accessing cached descendants before the requested path is updated (i.e., until the cache updates for a multi-path write are completed), so as to ensure cache consistency.
VI Local Hash Collision Resolution
@@
Figure 5: Example of token generation and distribution. Note that the controller also assigns tokens for $p$’s uncached ancestors in Steps 2-4. We omit the details for brevity.
@@
Fletch maps file-system pathnames to 64-bit hash keys (§IV-A), so hash collisions are possible (albeit unlikely) and lead to incorrect metadata retrieval. While the controller holds a global view of all cached paths (§IV-B), querying the controller to resolve hash collisions for every request is impractical due to high switch-to-controller latencies (§II-C). Fletch adopts a local, token-based hash collision resolution mechanism by synchronizing the controller’s global view of cached paths with the switch, clients, and servers, so as to ensure correctness without compromising performance.
@@
Hash collision resolution with tokens. A token is an 8-bit value paired with a 64-bit hash key to uniquely identify a path. Valid tokens range from 1 to 255, while 0 indicates invalid. A cached path is assigned token 1 if no collision occurs, or the next available token (e.g., 2) if it collides with an existing cached path. Thus, we can associate valid tokens with up to 255 paths with the same colliding hash key. For 64-bit hash keys, the collision probability between any two paths is $1/2^{64}$, making it unlikely for more than 255 paths to share the same hash key in a large namespace. Thus, 8-bit tokens suffice for hash collision resolution.
@@
Fletch maintains two unordered map structures: (i) a path-token map, which records paths and their tokens (path-token pairs), and (ii) a hash-token map, which records hash keys and their tokens (hash-token pairs). The controller keeps both maps, each client and server holds a path-token map, and the switch holds a hash-token map.
@@
Token generation and distribution. During cache admission, the controller assigns tokens to each level of a hot path, as shown in Figure 5. When the switch reports a hot path $p$ for admission ( 1), the controller checks its map structures. If $p$ is the first time being admitted and its hash key is unique, the controller assigns a token of value one; if a collision occurs, it assigns the next available token (e.g., $t$) ( 2). The controller updates its path-token and hash-token maps with the new path-token and hash-token pairs, respectively ( 3). These entries persist in the controller’s maps even after cache eviction to ensure consistency. If an evicted path is later re-admitted, its previously assigned token is reused. In the worst case, the controller may store entries for all paths ever cached. For scalability, the controller can use persistent key-value stores (e.g., RocksDB [rocksdb]) for the map structures.
@@
Figure 6: Example of how a client updates its path-token map. Note that the client also attaches tokens of $p$’s ancestors in Step 1, the server returns the tokens for $p$’s ancestors in Step 5, and the client adds the tokens in Step 6. We omit the details for brevity.
@@
Discussion. Fletch’s token-based mechanism ensures correctness with modest client complexity and overhead. Clients must maintain path-token maps and attach tokens to requests, and incur a cache miss on the first read of a cached path due to the initial lack of a valid token. Also, including tokens in requests adds minor communication overhead. Theoretically, the number of tokens grows linearly with the depth of a path. Nevertheless, as most paths have small depths (§V-A), the communication overhead is minimal. If the controller fails, Fletch needs to reboot the switch, controller, and servers to ensure correct cache retrieval. The current caches can be quickly recovered by exporting all caches before the reboot.
VII Implementation
We have built a prototype of Fletch, including the controller, in-switch cache, clients, and servers.
Controller. The controller is implemented in C++ with 2.8 K LoC. It leverages APIs provided by the Tofino switch compiler [tofino] to interact with the switch data plane, including configurations and updates of MATs and registers, for cache admission and eviction (§IV) and token management (§VI).
In-switch cache. The in-switch cache is implemented in P4 with 5.1 K LoC and compiled for a Tofino switch [tofino, switchthpt]. It comprises several components: (i) a hash-token map using MATs for cache lookups and local hash collision resolution, where each entry is a 9-byte key (8-byte hash key and 1-byte token); (ii) 32 register arrays of 32-bit slots for storing cache entries; (iii) a three-row CMS with three register arrays, each with 64K 16-bit slots; (iv) a frequency counter array with a register array of 32-bit slots; (v) lock counter arrays with eight register arrays, each with 16K 16-bit slots; (vi) a validation array with a register array of 1-bit slots; and (vii) a sequence counter array with a register array of 8-bit slots.
Client implementation. The C++ client driver supports multi-threaded workload execution, where each thread simulates a logical client. Each client maintains a path-token map for local hash collision resolution (§VI). It communicates with the switch via UDP-encapsulated packets customized for metadata requests, with re-transmission support for reliability. We integrate the driver with the mdtest [mdtest] benchmarking tool for HDFS metadata operations.
Server implementation. Each server is implemented in C++ with 3 K LoC and hosts an HDFS (v3.2) namenode [hdfs]. We use RBF [hdfsrbf] with the HASH_ALL policy, which uses consistent hashing to distribute files evenly across all namenodes for load balancing and creates directories on all namenodes. Each server manages a subset of the metadata namespace, connects to its local namenode via the C++ HDFS client library libhdfs3 [libhdfs3], and serves client requests while maintaining a path-token map for local hash resolution (§VI).
VIII Evaluation
VIII-A Methodology
Table I: Real-world workloads and their metadata operation ratios.
@@ VIII-A Methodology
Alibaba [lv22] 52.6% open/close, 9.59% create, 3.9% readdir, 69.1%
@@
Training [xu24] 54.32% open/close, 28.5% stat, 0.13% readdir, 83.1%
@@
Thumb [xu24] 57.01% open/close, 28.44% stat, 0.13% readdir, 85.7%
@@
LinkedIn [ren14] 84% open/getattr, 9% create/mkdir, 84%
@@
Testbed. We evaluate Fletch on a testbed comprising a 3.2 Tbps two-pipeline Tofino switch [tofino, switchthpt] and three physical machines. Two machines host servers, each with a 2.40 GHz 10-core Intel Xeon Silver 4210R CPU, 128 GiB DRAM, and a 2 TB HDD (Dell PERC H330 Mini). The client driver runs on the remaining machine, with a 2.40 GHz, 16-core Intel Xeon Silver 4314 CPU, 128 GiB DRAM, and a 960 GB NVMe SSD (Micron 9300 PRO). Each machine is connected to the switch via a 40 Gbps NIC (Mellanox ConnectX-5 CX516A). The client machine uses one pipeline, and the two server machines use another pipeline. All counter arrays (§VII) reside in the server-connected pipeline.
@@
Since our two-pipeline Tofino switch lacks native support for cross-pipeline recirculation, we physically connect the designated ingress pipeline hosting the lock counter arrays to another ingress pipeline using a physical wire [sheng25]. This enables requests from another ingress pipeline to be recirculated to the designated ingress pipeline (§V). Future programmable switches with native cross-pipeline recirculation would eliminate this requirement.
@@
Workloads. We evaluate Fletch using four real-world workload traces: (i) Alibaba’s Pangu file-system instances (Alibaba) [lv22], (ii) convolutional neural network training (Training) [xu24], (iii) the processing of one million thumbnail images (Thumb) [xu24], and (iv) a LinkedIn HDFS cluster (LinkedIn) [ren14]. Table I summarizes the proportion of metadata operations and the read ratio for each workload.
@@
We refine the workloads for our evaluation. To focus on metadata performance, we exclude reads and writes of file data, following prior studies [niazi17, li17, ren14, lv22]. For Training and Thumb that include file reads and writes [xu24], we exclude these operations and normalize the remaining metadata operations. Since we exclude file writes, we treat close as a read operation, while it updates both modification and access timestamps if the closed file has been updated. HDFS updates access timestamps hourly by default [defaultsetting], and we exclude timestamp updates as they are infrequent.
@@
For LinkedIn, as the original paper [ren14] does not provide operation ratios, we adjust the ratios based on trace analysis [lv22] as follows: open (42%), getattr (42%), create (4.5%), mkdir (4.5%), chmod (1%), delete (3%), and rename (3%). We assign the smallest ratio to chmod as it is less frequent than delete and rename [lv22]. Further, getattr corresponds to stat and statdir in HDFS. We replace getattr with stat, as file operations are much more frequent than directory operations [lv22], and focus on file renaming for rename.
@@
Generation of metadata operations. We use mdtest [mdtest] to generate file-system namespaces and metadata operations. By default, we configure a path depth of nine, as metadata requests are often aggregated at small depths [agrawal07, meyer12, douceur99]V-A). We create 32 million empty files to focus on metadata performance; empty files are also used in prior evaluation [ren14, xiao15, niazi17, li17, lv22, wang23]. To simulate workload skewness, we use the 80/20 rule, with 80% of operations on 20% of directories and files [wang21, xu24]. We model file access frequencies across levels using a power-law distribution with an exponent of 0.9 and randomly assigning an operation type using the ratios in Table I. For statdir and readdir, we choose the parent directory of the selected file. For mkdir and rmdir, we use separate directories to avoid removing non-empty directories. We address various workload settings in §VIII-C.
@@
We mix all metadata operations in each workload. However, rename, delete, and rmdir inherently involve metadata modifications that necessitate the granting and revoking of leases in HDFS [wang09]. The lease-based operations on frequently created and deleted files can slow down all metadata operations when we perform stress tests. Thus, we place rename, delete, and rmdir at the end of the request sequence, so as to keep metadata operations at high rate.